# Performances of a radial TPC for the detection of neutrinoless double beta decay

### v1

@@ 8.4 External neutrons
    Neutrons entering the detector can be captured by xenon, resulting in the emission of several gammas with a total energy of a few MeV. The expected neutron flux of $4\times 10^{-6}$ events cm-2 s-1 and the energy spectrum between 2 and 6 MeV were obtained from measurements performed at LSM NEMO:2002zps . Similar to the gamma simulation, neutrons were generated in the GEANT4 simulation on a sphere with a radius of 280 cm, homogeneously and with isotropic directions, yielding a statistics of about $1.2\times 10^{8}$ events per year. GEANT4 was also used to estimate neutron captures on the 136Xe isotope. The simulated statistics of $3\times 10^{10}$ events, corresponding to approximately 240 years of data taking, a the total number of background events after cuts of 0.02.
@@
The neutron related background expectation is summarized in Tab. 4
@@ 8.7 Background summary
Composite radioactivity 1.3
@@ 9 Pile Up
The total event rate is a critical parameter for the experiment. Given the long drift time, on the order of a few milliseconds, even events outside the ROI, though not directly contributing to the background, could overlap with good events, compromising their energy or position reconstruction. To maintain a minimal pile-up probability of $5\times 10^{-3}$, assuming a time window of 5 ms which corresponds to the maximum expected drift time, the event rate should be kept below 1 Hz. The same background sources considered in Sec.8 were re-evaluated with a significantly lower energy threshold, down to a few keV.
@@
The radioactivity of the composite material was simulated in Sec. 8.2, focusing on 60Co, 232Th, and 238U, with an assumed activity of 10 $\mu$Bq/kg. The fraction of events releasing energy within the xenon active volume is 43%, 8%, and 5%, respectively. Based on the expected number of events, the rates for these decay chains are $1.4\times 10^{-3}$ Hz, $2.2\times 10^{-3}$ Hz, and $2.2\times 10^{-3}$ Hz, yielding a total rate of $5.8\times 10^{-3}$ Hz.
@@
The total event rate expected in the detector is approximately 0.19 Hz (see Tab. 6), which is significantly below the acceptable limit of 1 Hz, considered necessary for safe detector operation.
@@
Composite radioactivity $5.8\times 10^{-3}$
@@ 10 Sensitivity results
where $G^{0\nu}$ represents the phase space factor and $M^{0\nu}$ is the matrix element. Using the values from Ref. KamLAND-Zen:2022tow and the references therein, the limit on $<m_{\beta\beta}>$ was computed for zero background, and the current estimate is shown in Fig. 13. Considering the expected background, the limit on $<m_{\beta\beta}>$ lies in the range of 7 meV to 33 meV after 5 years, which is reduced to 5 meV to 21 meV if data collection over 10 years is considered.
@@ 11 R2D2 in the worldwide panorama
Depending on the final background, the R2D2 proposed detector could completely rule out the inverted mass hierarchy region and begin to address part of the normal one. The performance of the proposed experiment can be compared with future multi-ton experiments. In particular, the following experiments were considered: nEXO nEXO:2021ujk , LEGEND1000 LEGEND:2021bnm ; LEGEND1000 , and CUPID CUPID:2022wpt . As shown in Fig. 14, the performance of R2D2 is comparable to that of future proposed experiments and could be even better if zero background is achieved. A reduction of the external gamma component (0.1 events per year) is relatively straightforward, as it depends solely on shielding, whereas the ultimate background arises from composite radioactivity (1.3 events per year) and radon contamination (0.2 events per year). As previously explained, a radioactivity level of 10 $\mu$Bq/kg is considered in the present study, but efforts are ongoing to potentially reduce this value.
Another point of comparison could be costs and timescale. The timescale is nearly identical for all next-generation experiments, which aim to yield final results around 2040. In this context, R2D2 could be timely, as final results could be expected around 2040 if the program is fully funded and data collection commences in 2030. The cost of the detector is exceptionally low, making R2D2 more than an order of magnitude cheaper than any other future proposed experiment. Furthermore, no cryogenic infrastructure is needed, and the power consumption is minimal, reducing the environmental impact.
@@
Figure 14: Expected limits for $<m_{\beta\beta}>$ at 90% C.L. for various experiments.

### v2

@@ 8.6 External neutrons
    Neutrons entering the detector can be captured by xenon, resulting in the emission of several gammas with a total energy of a few MeV. The expected neutron flux of $4\times 10^{-6}$ events cm-2 s-1 and the energy spectrum between 2 and 6 MeV were obtained from measurements performed at LSM NEMO:2002zps . Similar to the gamma simulation, neutrons were generated in the GEANT4 simulation on a sphere with a radius of 280 cm, homogeneously and with isotropic directions, yielding statistics of about $1.2\times 10^{8}$ events per year. GEANT4 was also used to estimate neutron captures on the 136Xe isotope. The simulated statistics of $3\times 10^{10}$ events, corresponding to approximately 240 years of data taking, a the total number of background events after cuts of 0.02.
@@
The neutron related background expectation is summarized in Tab. 4.
@@ 8.8 Cosmogenic background
A muon veto does not seems necessary given the weak production rate of 137Xe, however it could be considered if needed depending on the actual experiment location.
@@ 8.9 Background summary
Figure 15: Signal and background energy spectrum before 15(a) and after  15(b) selection cuts. The different components (thin lines) as well the sum of all the background (thick black line) are shown. For comparison a signal expected for a mean lifetime of $T^{0\nu}_{1/2}=10^{26}$ years is also shown (thick blue line).
@@
The background energy spectrum is also shown in Fig. 15 before and after the application of the selection cuts.
@@
Composite radioactivity 1.2
Anode radioactivity $1.0\times 10^{-3}$
Lead radioactivity 0.1
@@ 9 Pile Up
The total event rate is a critical parameter for the experiment. Given the long drift time, of the order of a few milliseconds, even events outside the ROI, though not directly contributing to the background, could overlap with good events, compromising their energy or position reconstruction. To maintain a minimal pile-up probability of $5\times 10^{-3}$, assuming a time window of 5 ms which corresponds to the maximum expected drift time, the event rate should be kept below 1 Hz. The same background sources considered in Sec. 8 were re-evaluated with a significantly lower energy threshold, down to a few keV.
@@
The radioactivity of the composite material was simulated in Sec. 8.2, focusing on 232Th and 238U, with an assumed activity of 10 $\mu$Bq/kg. The fraction of events releasing energy within the xenon active volume is 8% and 5%, respectively. Based on the expected number of events, the rate for each of these decay chains is $2.2\times 10^{-3}$ Hz, yielding a total rate of $4.4\times 10^{-3}$ Hz.
The activity of the anode has a negligible impact on pile-up, as its small mass results in an expected event rate of approximately 1500 events per year, corresponding to $4.7\times 10^{-5}$ Hz.
The activity of lead also has a negligible impact on pile-up, since even without any reduction, as discussed in Sec. 8.4, the contribution to the pile-up is at the level of $8.0\times 10^{-3}$ Hz.
@@
The total event rate expected in the detector is approximately 0.20 Hz (see Tab. 6), which is significantly below the acceptable limit of 1 Hz, considered necessary for safe detector operation.
@@
Composite radioactivity $4.4\times 10^{-3}$
Anode radioactivity $4.7\times 10^{-5}$
Lead radioactivity $8.0\times 10^{-3}$
@@ 10 Sensitivity results
where $G^{0\nu}$ represents the phase space factor and $M^{0\nu}$ is the matrix element. Using the values from Ref. KamLAND-Zen:2022tow and the references therein, the limit on $<m_{\beta\beta}>$ was computed both for the zero-background scenario and for the current background estimate, with the results shown in Fig. 17. Considering the expected background, the limit on $<m_{\beta\beta}>$ lies in the range of 13 meV to 57 meV after 10 years, which is reduced to 8 meV to 35 meV if zero background is achieved.
@@ 11 R2D2 in the worldwide panorama
The present best senstivities on the effective neutrino mass have been obtained by KamLAND-Zen KamLAND-Zen:2024eml and GERDA GERDA:2020xhi . KamLAND-Zen measured $2.097~{}\mathrm{ton\cdot yr}$ of 136Xe, achieving a $\beta\beta 0\nu$ half-life limit of $T_{1/2}^{0\nu}>3.8\times 10^{26}~{}\mathrm{years}$ at 90% C.L., corresponding to an effective neutrino mass $<m_{\beta\beta}>$ between 28 and 122 meV. GERDA measured $125~{}\mathrm{kg\cdot yr}$ of 76Ge, setting a half-life limit of $T_{1/2}^{0\nu}>1.8\times 10^{26}~{}\mathrm{years}$, which corresponds to $<m_{\beta\beta}>$ between 79 and 180 meV.
The performance of the proposed experiment depends on the final background and could be compared with future multi-ton experiments. In particular, the following experiments were considered: nEXO nEXO:2021ujk , LEGEND1000 LEGEND:2021bnm ; LEGEND1000 , and CUPID CUPID:2022wpt .
As shown in Fig. 18, the performance of R2D2 is only a factor of 2.5 worse with respect to that of future proposed experiments and could be comparable if zero background is achieved. A reduction of the external gamma component (0.1 events per year) is relatively straightforward, as it depends solely on shielding, whereas the ultimate background arises from composite radioactivity (1.2 events per year) and radon contamination (0.2 events per year). As previously explained, a radioactivity level of 10 $\mu$Bq/kg is considered in the present study, but efforts are ongoing to potentially reduce this value. Furthermore, ongoing studies indicate a possible reduction in the thickness of the composite vessel by a factor of 6, resulting in a total vessel mass of only about 60 kg. Multiple detectors could also be built, increasing the active mass. The projected sensitivity for a 2 ton detector is also shown in Fig. 18 for comparison.
Another point of comparison could be costs and timescale. The timescale is nearly identical for all next-generation experiments, which aim to yield final results around 2040. In this context, R2D2 could be timely, as final results could be expected around 2040 if the program is fully funded and data collection commences in 2030. The cost of the detector itself is low ($\sim 3$ M€), excluding the cost of enriched xenon. Moreover, no cryogenic infrastructure is required, and the power consumption is minimal, which helps reduce the environmental impact.
@@
Figure 18: Current and expected limits for $<m_{\beta\beta}>$ at 90% C.L. for various experiments.